OpenStudy (anonymous):

Write the standard form of the line that passes through the given points. (7, -3) and (4, -8)

OpenStudy (amorfide):

Gave you a detailed response in this question before, try to work it out, we are not here to give you answers, only to help you, if you won't accept help, that is your problem

OpenStudy (anonymous):

I do not need your negativity. You did not explain it the correct way. Please do not comment again. I will find help from somebody else.

OpenStudy (anonymous):

Use the point slope formula: y – y1 = m(x – x1) m is the slope... slope is change in y divided by change in x then plug in a given point into y1 and x1.. you will then have the equation of the line in point slope form... after that rearrange the equation into standard form which is y = mx+b

OpenStudy (amorfide):

y=mx+c is the standard equation of a straight line to get this you must work out the y intercept, and the gradient the gradient of a line passing through two coordinates is the difference in y ordinates divided by the difference in x coordinates labelling the first set of coordinates to be (x1, y1) and the second set (x2, y2) the gradient would be equal to (y2-y1)/(x2-x1) so m=(y2-y1)/(x2-x1) next to get the equation of a line you use the formula y-y1=m(x-x1) where y1 is the y ordinate of one of the two coordinate sets, which ever one you choose, you must use the same set for the x ordinate for example if i did y-(-3) i would have to use 7 for the x1 to be consistent can you work out the answer? I will confirm if you are correct, or help along the way